# On homological properties of the Schlichting completion

### v1

@@ 2. The Schlichting completion
is the Schlichting completion of $\Gamma$ relative to the commensurated subgroup $\Lambda$.
In the following, we collect some properties of this completion.
@@
    The map $\alpha\colon\Gamma\to G$ is a dense embedding.
@@ 3. Finiteness properties of the Schlichting completion
If $\Lambda$ has type $\operatorname{FP}^{R}_{n}$ then so does any subgroup of $\Gamma$ that is commensurated with $\Lambda$ by [brown]*(5.1) Proposition on p. 197. Let $\Lambda^{\prime}<\Gamma$ commensurated with $\Lambda$. Let $Q_{\ast}\to R$ be a projective $R[\Lambda^{\prime}]$-resolution of the trivial module such that $Q_{0},\dots,Q_{n}$ are finitely generated. Then $R[\Gamma]\otimes_{R[\Lambda^{\prime}]}Q_{\ast}$ is a projective resolution of $R[\Gamma/\Lambda^{\prime}]$ that is finitely generated in degrees $0,\dots,n$. Hence the $R[\Gamma]$-module $R[\Gamma/\Lambda^{\prime}]$ has type $\operatorname{FP}^{R}_{n}$. A finitely generated discrete permutation $G$-module $R[\Omega]$ is a finite sum of modules of the type $R[G/U]$ where $U<G$ is a compact-open subgroup. By Lemma 2.2 we have $G/U\cong\Gamma/(\Gamma\cap U)$, and $\Gamma\cap U$ is commensurable with $\Gamma\cap\overline{\Lambda}=\Lambda$. Therefore $\operatorname{res}^{G}_{\Gamma}R[G/U]$ is of type $\operatorname{FP}^{R}_{n}$.
@@ 5. The Euler characteristic of the Schlichting completion
Lemma 5.1.
Let $\Gamma$ be a dense subgroup of an unimodular tdlc group $G$ with Haar measure $\mu$. Let $U$ and $V$ be compact-open subgroups of $G$. Assume that $\Gamma\cap U$ is of type $\operatorname{FP}^{\mathbb{Q}}$. Then
$\chi^{(2)}(\Gamma\cap U)=\frac{\mu(V)}{\mu(U)}\chi^{(2)}(\Gamma\cap V).$
Proof.
The groups $\Gamma\cap W$ where $W$ is any compact-open subgroup, are commensurable. Therefore, if one is of type $\operatorname{FP}^{\mathbb{Q}}$, then all are. This follows from the combination of [bieri]*Theorem 5.11 on p. 78 and [brown]*Proposition 5.1 on p. 197 and Proposition 6.1 on p. 199.
In particular, all these groups have a well-defined $\ell^{2}$-Euler characteristic. By applying the result for the pairs $U,V\cap U$ and $V,V\cap U$, we may assume that $V$ is a subgroup of $U$. By denseness of $\Gamma$ and discreteness of the target, the obvious map $(\Gamma\cap U)/(\Gamma\cap V)\to U/V$ is a bijection. Therefore
$[\Gamma\cap U:\Gamma\cap V]=\frac{\mu(U)}{\mu(V)}.$
Now the statement follows from (8). ∎
@@
Consider a projective resolution of the trivial $G$-module by proper discrete permutation modules as in (4). For every $j\in\{0,\dots,n\}$ and every $i\in I_{j}$ we choose a finite projective ${\mathbb{C}}[\Gamma\cap U_{i}^{(j)}]$-resolution $\tilde{Q}(i,j)_{\ast}$ of the trivial ${\mathbb{C}}[\Gamma\cap U_{i}^{(j)}]$-module ${\mathbb{C}}$. Since $\Gamma\cap U_{i}^{(j)}$ and $\Lambda$ are commensurable, the group $\Gamma\cap U_{i}^{(j)}$ is of type $\operatorname{FP}^{\mathbb{Q}}$ (thus, $\operatorname{FP}^{\mathbb{C}}$). See the proof of Lemma 5.1 for a reference. Tensoring this resolution with ${\mathbb{C}}[\Gamma]$ we obtain a finite projective ${\mathbb{C}}[\Gamma]$-resolution of ${\mathbb{C}}[\Gamma/(\Gamma\cap U_{i}^{(j)})]$ which we denote by $Q(i,j)_{\ast}$. For every $j\in\{0,\dots,n\}$, the sum $\bigoplus_{i\in I_{j}}Q(i,j)_{\ast}$ is a finite projective ${\mathbb{C}}[\Gamma]$-resolution of
$\operatorname{res}_{\Gamma}^{G}\bigl{(}\bigoplus_{i\in I_{j}}{\mathbb{C}}\bigl{[}G/U_{i}^{(j)}\bigr{]}\bigr{)}\cong\bigoplus_{i\in I_{j}}{\mathbb{C}}\bigl{[}\Gamma/(\Gamma\cap U_{i}^{(j)})\bigr{]}.$
@@
$=\sum_{j\geq 0}(-1)^{j}\sum_{i\in I_{j}}\sum_{k\geq 0}(-1)^{k}\dim_{L(\Gamma\cap U_{i}^{(j)})}\bigl{(}L(\Gamma\cap U_{i}^{(j)})\otimes_{{\mathbb{C}}[\Gamma\cap U_{i}^{(j)}]}\tilde{Q}(i,j)_{k}\bigr{)}$
$=\sum_{j\geq 0}(-1)^{j}\sum_{i\in I_{j}}\chi^{(2)}\bigl{(}\Gamma\cap U_{i}^{(j)}\bigr{)}$

### v2

@@ 2. The Schlichting completion
is the Schlichting completion of $\Gamma$ relative to the commensurated subgroup $\Lambda$. Strictly speaking, the Schlichting completion is not a completion of $\Gamma$ since $\alpha$ might not be injective.
In the following, we collect some properties of this construction.
@@
    The map $\alpha\colon\Gamma\to G$ has a dense image. Its kernel is the largest subgroup that is normal in $\Gamma$ and contained in $\Lambda$.
@@
An easy consequence is that $\alpha$ induces an isomorphism $\Gamma/\Lambda\xrightarrow{\cong}G/U$ where $G=\Gamma/\!/\Lambda$ and $U=\overline{\alpha(\Lambda)}$. This fact will be used frequently.
@@ 3. Finiteness properties of the Schlichting completion
If $\Lambda$ has type $\operatorname{FP}^{R}_{n}$ then so does any subgroup of $\Gamma$ that is commensurated with $\Lambda$ by [brown]*(5.1) Proposition on p. 197. Let $\Lambda^{\prime}<\Gamma$ commensurated with $\Lambda$. Let $Q_{\ast}\to R$ be a projective $R[\Lambda^{\prime}]$-resolution of the trivial module such that $Q_{0},\dots,Q_{n}$ are finitely generated. Then $R[\Gamma]\otimes_{R[\Lambda^{\prime}]}Q_{\ast}$ is a projective resolution of $R[\Gamma/\Lambda^{\prime}]$ that is finitely generated in degrees $0,\dots,n$. Hence the $R[\Gamma]$-module $R[\Gamma/\Lambda^{\prime}]$ has type $\operatorname{FP}^{R}_{n}$. A finitely generated discrete permutation $G$-module $R[\Omega]$ is a finite sum of modules of the type $R[G/U]$ where $U<G$ is a compact-open subgroup. By Lemma 2.2 we have $G/U\cong\Gamma/\alpha^{-1}(U)$, and $\alpha^{-1}(U)$ is commensurable with $\alpha^{-1}(\overline{\alpha(\Lambda)})=\Lambda$. Therefore $\operatorname{res}^{G}_{\Gamma}R[G/U]$ is of type $\operatorname{FP}^{R}_{n}$.
@@ 5. The Euler characteristic of the Schlichting completion
As before, we denote the canonical map of $\Gamma$ into the Schlichting completion $G=\Gamma/\!/\Lambda$ by $\alpha$. Further, $U$ is the closure of $\alpha(\Lambda)$. Consider a projective resolution of the trivial $G$-module by proper discrete permutation modules as in (4). For every $j\in\{0,\dots,n\}$ and every $i\in I_{j}$ we choose a finite projective ${\mathbb{C}}[\alpha^{-1}(U_{i}^{(j)})]$-resolution $\tilde{Q}(i,j)_{\ast}$ of the trivial ${\mathbb{C}}[\alpha^{-1}(U_{i}^{(j)})]$-module ${\mathbb{C}}$. Since $\alpha^{-1}(U_{i}^{(j)})$ and $\Lambda=\alpha^{-1}(U)$ are commensurable, the group $\alpha^{-1}(U_{i}^{(j)})$ is of type $\operatorname{FP}^{\mathbb{Q}}$ (thus, $\operatorname{FP}^{\mathbb{C}}$). This follows from the combination of [bieri]*Theorem 5.11 on p. 78 and [brown]*Proposition 5.1 on p. 197 and Proposition 6.1 on p. 199. Tensoring this resolution with ${\mathbb{C}}[\Gamma]$ we obtain a finite projective ${\mathbb{C}}[\Gamma]$-resolution of ${\mathbb{C}}[\Gamma/\alpha^{-1}(U_{i}^{(j)})]$ which we denote by $Q(i,j)_{\ast}$. For every $j\in\{0,\dots,n\}$, the sum $\bigoplus_{i\in I_{j}}Q(i,j)_{\ast}$ is a finite projective ${\mathbb{C}}[\Gamma]$-resolution of
$\operatorname{res}_{\Gamma}^{G}\bigl{(}\bigoplus_{i\in I_{j}}{\mathbb{C}}\bigl{[}G/U_{i}^{(j)}\bigr{]}\bigr{)}\cong\bigoplus_{i\in I_{j}}{\mathbb{C}}\bigl{[}\Gamma/\alpha^{-1}(U_{i}^{(j)})\bigr{]}.$
@@
$=\sum_{j\geq 0}(-1)^{j}\sum_{i\in I_{j}}\sum_{k\geq 0}(-1)^{k}\dim_{L(\alpha^{-1}(U_{i}^{(j)}))}\bigl{(}L(\alpha^{-1}(U_{i}^{(j)}))\otimes_{{\mathbb{C}}[\alpha^{-1}(U_{i}^{(j)})]}\tilde{Q}(i,j)_{k}\bigr{)}$
$=\sum_{j\geq 0}(-1)^{j}\sum_{i\in I_{j}}\chi^{(2)}\bigl{(}\alpha^{-1}(U_{i}^{(j)})\bigr{)}$